### 18 Dec 2023 v4

@@ I Introduction
Due to the rise of social media, tens of millions of images are generated and transmitted on the web every second. In order to conserve bandwidth, service providers are compelled to seek more efficient and effective image compression methods. Although traditional coding methods like JPEG [1], JPEG2000 [2], AVC [3], HEVC [4, 5], and VVC [6] have achieved commendable performance, their design relies on manual design for each module. This lack of joint optimization hampers their ability to fully exploit the potential for further advancements in image compression.
Recently, various learned image compression models  [7, 8, 9, 10, 11, 12, 13] have emerged, showcasing impressive performance gains.
@@
State-of-the-art learned image compression models [15, 16, 20, 21, 31, 22, 27, 29] commonly enhance the entropy model by incorporating a hyper-prior module [9] or a context module [11]. These additional modules enable the estimation of conditional entropy and the utilization of conditional probabilities for entropy coding. Context modules usually model probabilities and correlations in different dimensions, including local spatial context module, global spatial context module, and channel-wise context module. However, the current global context modules rely on computationally intensive quadratic complexity computations, which consumes huge GPU memories and have slower encoding and decoding speed, imposing limitations on the potential for high-resolution image coding. Furthermore, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To overcome the aforementioned limitations, we propose a novel linear complexity attention-based multi-reference entropy model. This entropy model effectively captures local spatial, global spatial, and channel-wise contexts with linear complexity and can be employed for efficient high-resolution image coding, which is denoted as ${}^{++}$ to differ from our prior work [27] presented at ACMMM 2023. Based on ${}^{++}$, we introduce ${}^{++}$, which achieve state-of-the-art performance.
@@ II-A Learned Image Compression
where $\bm{x}$ represents the input image, $g_{a}$ transform the $\bm{x}$ to compact latent representation $\bm{y}$. $\bm{y}$ is quantized to $\hat{\bm{y}}$ for entropy coding. $\hat{\bm{x}}$ represents the decompressed image. $\theta$ and $\phi$ are parameters of $g_{a}$ and $g_{s}$. Since quantization is non-differentiable, it can be addressed during training by either adding uniform noise $\mathcal{U}(-0.5,0.5)$ [7, 9] or using the straight-through estimator (STE) [8]. In particular, when uniform noise is added, the rate-distortion optimization target in Equation 1 is equivalent to evidence lower bound (ELBO) optimization in variational auto-encoders [30]. To enhance non-linearity, Generalized Divisive Normalization (GDN) [36] layers or its variants [23, 37] are employed. Additionally, self-attention [38, 39, 40, 21, 41], ensemble techniques [42], and block partition [19] are utilized in transform modules for more compact latent representations. In the basic model, a factorized or a non-adaptive density entropy model is adopted.
@@ III-E Linear Complexity Intra-Slice Global Context Module
In latent representation $\hat{\bm{y}}\in\mathbb{R}^{L\times C}$, where $L=H\times W$, $H$ is the height of latent representation, $W$ is the width of latent representation, $C$ is the channel number, each channel contains distinct information, but they can be considered as thumbnails, as depicted in Fig. 2. Notably, the channels exhibit similar global similarities. This is evident from the visualization of cosine similarities between two slices of Cheng’20 [15], as visualized in Fig. 8, where despite differences in magnitude, the global correlations are similar. When decoding the current slice $\hat{\bm{y}}^{i}\in\mathbb{R}^{L\times S}$, decoded slice $\hat{\bm{y}}^{i-1}\in\mathbb{R}^{L\times S}$ assists in estimating the global correlations in slice $\hat{\bm{y}}^{i}$. However, a challenge arises in determining how to estimate these global correlations. While cosine similarity may be useful, it is fixed and may not accurately capture the features. In this regard, attention maps prove to be a suitable choice. The embedding layer is learnable, which make it flexible to adjust the method for global correlations estimation by modifying queries, keys, and values.
@@ IV-B Benchmarks and Metrics
  • Kodak [52] is selected as a test set for almost all end-to-end image compression models [7, 8, 9, 11, 15, 16, 17, 18, 19, 20, 24, 25, 21, 22, 26, 27, 28, 29, 41]. It contains 24 raw $768\times 512$ images.
@@
  • CLIC Professional Valid is the validation set of 3rd Challenge on Learned Image Compression which contains $41$ images. Image in this dataset contains around $2048\times 1440$ pixels. CLIC Pro Valid is widely used in many recent methods [15, 20, 22, 21, 41, 27, 54, 55, 28].
@@ IV-C1 Quantitive Results
BD-rate reductions are presented in Table III. When computing BD-rate, VTM-17.0 Intra under YUV444 is employed as anchor. We use unofficial weights of ELIC [^2] https://github.com/VincentChandelier/ELiC-ReImplemetation to evaluate the rate-distortion performance of ELIC on Tecnick [53], CLIC Professional Valid [14]. Our ${}^{++}$ outperforms our previous MLIC and ${}^{+}$ [27]. Our ${}^{++}$ reduces BD-rate by $13.39\%$ over VTM-17.0 Intra on Kodak [52] while existing method ELIC [22] only reduces $5.95\%$ BD-rate and LIC-TCM [41] only reduces $10.14\%$ BD-rate on Kodak. Moreover, our ${}^{++}$ performs much better on high-resolution datasets, such as Tecnick [53] and CLIC Professional Valid [14]. Our ${}^{++}$ reduces $17.59\%$ on Tecnick [53] and reduces $13.08\%$ BD-rate on CLIC Professional Valid [14], which are much better than existing methods. Our ${}^{++}$ achieves state-of-the-art performance on all three datasets. The excellent performance of our ${}^{++}$ on these datasets also demonstrates the excellent generalization of our ${}^{++}$. We highlight the BD-rate for MS-SSIM [58], our proposed ${}^{++}$ reduces about $50\%$ bits compared to VTM-17.0 Intra, which is a large progress in learned image compression.
@@ IV-C2 Qualitive Results
To further demonstrate superiority of our proposed ${}^{++}$, we compare our ${}^{++}$ with learned image compression models Cheng’20 [15], Xie’21 [20], STF [21], WACNN [21], ELIC [22] and non-neural codec VTM-17.0 Intra [6] on perceptual quality. Fig. 11 presents the reconstructions of Kodim07 from Kodak. PSNR value of the image reconstructed by our ${}^{++}$ optimized for MSE is $1$dB higher than image reconstructed by VTM-17.0 Intra [6]. MS-SSIM [58] of the image reconstructed by our ${}^{++}$ optimized for MS-SSIM is $0.02$ higher than image reconstructed by VTM-17.0 Intra. The windowsill of reconstructions are cropped to patches for clearer comparisons. The reconstructions of ${}^{++}$ have sharper textures retain more details. In terms of visual quality, our ${}^{++}$ have significant improvements on rate-perception performance compared to other models.
@@ IV-D Computational Complexity
In order to better fit the real-world scenario, we test the model complexities in the case of different resolution images as inputs. We select $16$ images with resolution larger than $3584\times 3584$ from LIU4K test dataset [66] and we center crop these images to $\{512\times 512,768\times 768,1024\times 1024,1536\times 1536,2048\times 2048,2560\times 2560,3072\times 3072,3584\times 3584\}$ patches. We compared our ${}^{++}$ with our prior work ${}^{+}$ [27] which employs vanilla attention to illustrate the advantages of proposed linear complexity global context capturing. We also compare our proposed ${}^{++}$ with recent learned image compression models [15, 23, 20, 34, 54, 21, 22, 41]. The experiments are conducted on a Tesla A100 GPU and a Xeon(R) Platinum 8260C CPU. The results are illustrated in Fig. 12.
@@ IV-D1 On GPU Memory comsuption
The quadratic complexity of vanilla attention leads to significantly more memory consumptions on high resolution image coding. When compressing $2048\times 2048$ images, ${}^{+}$ consumes nearly $22.37$ GB GPU memory. Our proposed linear complexity global context capturing significantly reduces the consumption of GPU memory as our proposed ${}^{++}$ only takes about $5.55$ GB GPU memory to compress a $2048\times 2048$ image. When compressing $3072\times 3072$ images, ${}^{+}$ consumes $46.38$ GB GPU memory while our proposed ${}^{++}$ only consumes $7.69$ GB GPU memory. Compared with recent LIC-TCM [41], our ${}^{++}$ consumes $\frac{1}{2}$ of the GPU memory consumed by LIC-TCM [41] when compressing $2048\times 2048$ images. When compressing $2560\times 2560$ images, our ${}^{++}$ only consumes about $\frac{1}{4}$ of the GPU memory consumed by LIC-TCM [41]. When compressing a $3584\times 3584$ image, the GPU memory consumption of our proposed ${}^{++}$ is still only $12.3$ GB while the LIC-TCM [41] requires $45.95$ GB GPU memory. The curve of the GPU memory consumed by our ${}^{++}$ as the resolution grows is also much flatter.
@@ IV-D3 On Forward MACs
Compared with ${}^{+}$, our ${}^{++}$, which employs the proposed linear complexity global spatial context modules, has lower MACs. Compared with the recent LIC-TCM [41], our ${}^{++}$ demonstrates significantly reduced MACs. Compared with ELIC [22], STF [21], and WACNN [21], our ${}^{++}$ has higher MACs when the input is high-resolution image. One contributing factor to the higher MACs is the utilization of Cheng’20 [15] as the basis for the transform module in our proposed ${}^{++}$. While the context module in Cheng’20 [15] is simpler than that of ELIC [22], STF [21], and WACNN [21], the overall MACs of Cheng’20 [15] are higher due to its transform modules having higher MACs. Since our ${}^{++}$ employs a modification of analysis transform and synthesis transform of Cheng’20 [15], it results in higher MACs compared to ELIC [22] and WACNN [21]. In addition, since our ${}^{++}$ employed an more advanced entropy model, which leads to higher MACs. However, the entropy model has modest effect on the overall MACs since the input image is down-sampled for four times in analysis transform, which implies that designing more advanced entropy models is more resource-efficient.
@@ IV-E5 Analysis of Inter-Slice Global Context Module
TABLE V: Influences of Patch Size on Rate-Distortion Performance. BD-rate (%) is employed as the metric. The anchor is ${}^{++}$ trained on $256\times 256$ patches.
@@
TABLE VI: Parameter number and Forward Macs of each context modules on Kodak dataset.

### 07 Jan 2024 v5

@@ I Introduction
Due to the rise of social media, tens of millions of images are generated and transmitted on the web every second. In order to conserve bandwidth, service providers are compelled to seek more efficient and effective image compression methods. Although traditional coding methods like JPEG [1], JPEG2000 [2], AVC [3], HEVC [4], and VVC [5] have achieved commendable performance, their design relies on manual design for each module. This lack of joint optimization hampers their ability to fully exploit the potential for further advancements in image compression.
Recently, various learned image compression models  [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] have emerged, showcasing impressive performance gains.
@@
State-of-the-art learned image compression models [24, 23, 28, 38, 29, 30, 36] commonly enhance the entropy model by incorporating a hyper-prior module [17] or a context module [19]. These additional modules enable the estimation of conditional entropy and the utilization of conditional probabilities for entropy coding. Context modules usually model probabilities and correlations in different dimensions, including local spatial context module, global spatial context module, and channel-wise context module. However, the current global context modules rely on computationally intensive quadratic complexity computations, which consumes huge GPU memories and have slower encoding and decoding speed, imposing limitations on the potential for high-resolution image coding. Furthermore, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To overcome the aforementioned limitations, we propose a novel linear complexity attention-based multi-reference entropy model. This entropy model effectively captures local spatial, global spatial, and channel-wise contexts with linear complexity and can be employed for efficient high-resolution image coding, which is denoted as ${}^{++}$ to differ from our prior work [30] presented at ACMMM 2023. Based on ${}^{++}$, we introduce ${}^{++}$, which achieve state-of-the-art performance.
@@ II-A Learned Image Compression
where $\bm{x}$ represents the input image, $g_{a}$ transform the $\bm{x}$ to compact latent representation $\bm{y}$. $\bm{y}$ is quantized to $\hat{\bm{y}}$ for entropy coding. $\hat{\bm{x}}$ represents the decompressed image. $\theta$ and $\phi$ are parameters of $g_{a}$ and $g_{s}$. Since quantization is non-differentiable, it can be addressed during training by either adding uniform noise $\mathcal{U}(-0.5,0.5)$ [15, 17] or using the straight-through estimator (STE) [16]. In particular, when uniform noise is added, the rate-distortion optimization target in Equation 1 is equivalent to evidence lower bound (ELBO) optimization in variational auto-encoders [37]. To enhance non-linearity, Generalized Divisive Normalization (GDN) [42] layers or its variants [26, 8] are employed. Additionally, self-attention [43, 44, 45, 25, 27, 10, 14], ensemble techniques [9], and block partition [6] are utilized in transform modules for more compact latent representations. In the basic model, a factorized or a non-adaptive density entropy model is adopted.
@@ III-E Linear Complexity Intra-Slice Global Context Module
In latent representation $\hat{\bm{y}}\in\mathbb{R}^{L\times C}$, where $L=H\times W$, $C$ is the channel number, each channel contains distinct information, but they can be considered as thumbnails, as depicted in Fig. 2. Notably, the channels exhibit similar global similarities. This is evident from the visualization of cosine similarities between two slices of Cheng’20 [24], as visualized in Fig. 8, where despite differences in magnitude, the global correlations are similar. When decoding the current slice $\hat{\bm{y}}^{i}\in\mathbb{R}^{L\times S}$, decoded slice $\hat{\bm{y}}^{i-1}\in\mathbb{R}^{L\times S}$ assists in estimating the global correlations in slice $\hat{\bm{y}}^{i}$. However, a challenge arises in determining how to estimate these global correlations. While cosine similarity may be useful, it is fixed and may not accurately capture the features. In this regard, attention maps prove to be a suitable choice. The embedding layer is learnable, which make it flexible to adjust the method for global correlations estimation by modifying queries, keys, and values.
@@ IV-B Benchmarks and Metrics
  • Kodak [55] is selected as a test set for almost all end-to-end image compression models [15, 16, 17, 19, 24, 23, 31, 14, 6, 28, 32, 33, 25, 29, 34, 30, 35, 27]. It contains 24 raw $768\times 512$ images.
@@
  • CLIC Professional Valid [22] is the validation set of 3rd Challenge on Learned Image Compression which contains $41$ images. Image in this dataset contains around $2048\times 1440$ pixels. CLIC Pro Valid is widely used in many recent methods [24, 28, 29, 25, 27, 30, 57, 58, 35].
@@ IV-C1 Quantitive Results
BD-rate reductions are presented in Table III. When computing BD-rate, VTM-17.0 Intra under YUV444 is employed as anchor. Unofficial weights of ELIC [^2] https://github.com/VincentChandelier/ELiC-ReImplemetation are used to evaluate the rate-distortion performance of ELIC on Tecnick, CLIC Professional Valid. Our ${}^{++}$ outperforms our previous MLIC and ${}^{+}$ [30]. Our ${}^{++}$ reduces BD-rate by $13.39\%$ over VTM-17.0 Intra on Kodak while existing method ELIC only reduces $5.95\%$ BD-rate and LIC-TCM only reduces $10.14\%$ BD-rate on Kodak. Moreover, our ${}^{++}$ performs better on high-resolution datasets, such as Tecnick and CLIC Professional Valid. Our ${}^{++}$ reduces $17.59\%$ on Tecnick and reduces $13.08\%$ BD-rate on CLIC Professional Valid, which are much better than existing methods. Our ${}^{++}$ achieves state-of-the-art performance on all three datasets. The excellent performance of our ${}^{++}$ on these datasets also demonstrates the excellent generalization of our ${}^{++}$. We highlight the BD-rate for MS-SSIM, our proposed ${}^{++}$ reduces about $50\%$ bits compared to VTM-17.0 Intra, which is a large progress in learned image compression.
@@ IV-C2 Qualitive Results
To further demonstrate superiority of our proposed ${}^{++}$, we compare our ${}^{++}$ with learned image compression models Cheng’20 [24], Xie’21 [28], STF [25], WACNN [25], ELIC [29] and non-neural codec VTM-17.0 Intra [5] on perceptual quality. Fig. 11 presents the reconstructions of Kodim07 from Kodak. PSNR value of the image reconstructed by our ${}^{++}$ optimized for MSE is $1$dB higher than image reconstructed by VTM-17.0 Intra [5]. MS-SSIM of the image reconstructed by our ${}^{++}$ optimized for MS-SSIM is $0.02$ higher than image reconstructed by VTM-17.0 Intra. The windowsill of reconstructions are cropped to patches for clearer comparisons. The reconstructions of ${}^{++}$ have sharper textures retain more details. In terms of visual quality, our ${}^{++}$ have significant improvements on rate-perception performance compared to other models.
@@ IV-D Computational Complexity
In order to better fit the real-world scenario, we test the model complexities in the case of different resolution images as inputs. We select $16$ images with resolution larger than $3584\times 3584$ from LIU4K test dataset [69] and we center crop these images to $\{512\times 512,768\times 768,1024\times 1024,1536\times 1536,2048\times 2048,2560\times 2560,3072\times 3072,3584\times 3584\}$ patches. We compared our ${}^{++}$ with our prior work ${}^{+}$ which employs vanilla attention to illustrate the advantages of proposed linear complexity global context capturing. We also compare our proposed ${}^{++}$ with recent learned image compression models [24, 26, 28, 41, 57, 25, 29, 27]. The experiments are conducted on a Tesla A100 GPU and a Xeon(R) Platinum 8260C CPU. The results are illustrated in Fig. 12.
@@ IV-D1 On GPU Memory comsuption
The quadratic complexity of vanilla attention leads to significantly more memory consumptions on high resolution image coding. When compressing $2048\times 2048$ images, ${}^{+}$ consumes nearly $22.37$ GB GPU memory. Our proposed linear complexity global context capturing significantly reduces the consumption of GPU memory as our proposed ${}^{++}$ only takes about $5.55$ GB GPU memory to compress a $2048\times 2048$ image. When compressing $3072\times 3072$ images, ${}^{+}$ consumes $46.38$ GB GPU memory while our proposed ${}^{++}$ only consumes $7.69$ GB GPU memory. Compared with recent LIC-TCM [27], our ${}^{++}$ consumes $\frac{1}{2}$ of the GPU memory consumed by LIC-TCM when compressing $2048\times 2048$ images. When compressing $2560\times 2560$ images, our ${}^{++}$ only consumes about $\frac{1}{4}$ of the GPU memory consumed by LIC-TCM. When compressing a $3584\times 3584$ image, the GPU memory consumption of our proposed ${}^{++}$ is still only $12.3$ GB while the LIC-TCM requires $45.95$ GB GPU memory. The curve of the GPU memory consumed by our ${}^{++}$ as the resolution grows is also much flatter.
@@ IV-D3 On Forward MACs
Compared with ${}^{+}$, our ${}^{++}$, which employs the proposed linear complexity global spatial context modules, has lower MACs. Compared with the recent LIC-TCM [27], our ${}^{++}$ demonstrates significantly reduced MACs. Compared with ELIC [29], STF [25], and WACNN [25], our ${}^{++}$ has higher MACs when the input is high-resolution image. One contributing factor to the higher MACs is the utilization of Cheng’20 [24] as the basis for the transform module in our proposed ${}^{++}$. While the context module in Cheng’20 is simpler than that of ELIC, STF, and WACNN, the overall MACs of Cheng’20 are higher due to its transform modules having higher MACs. Since our ${}^{++}$ employs a modification of analysis transform and synthesis transform of Cheng’20, it results in higher MACs compared to ELIC and WACNN. In addition, since our ${}^{++}$ employed an more advanced entropy model, which leads to higher MACs. However, the entropy model has modest effect on the overall MACs since the input image is down-sampled for four times in analysis transform, which implies that designing more advanced entropy models is more resource-efficient.
@@ IV-E5 Analysis of Inter-Slice Global Context Module
TABLE V: BD-RATE(%) of training using different patch size. The anchor is ${}^{++}$ trained on $256\times 256$ patches.
@@
TABLE VI: Parameters and Forward Macs of context modules on Kodak.